# DataEnvGym: Data Generation Agents in Teacher Environments with Student Feedback

### v1

@@ Abstract
The process of creating training data to teach models is currently driven by humans, who manually analyze model weaknesses and plan how to create data that improves a student model. Recent approaches using large language models (LLMs) as annotators reduce human annotation effort, but still require humans to interpret feedback from evaluations and control the LLM to produce data the student needs. Automating this labor-intensive process by creating autonomous data generation agents – or teachers – is desirable, but requires environments that can simulate the feedback-driven, iterative, closed loop of data creation. To enable rapid and scalable testing for such agents and their modules, we introduce DataEnvGym, a testbed of teacher environments for data generation agents. DataEnvGym frames data generation as a sequential decision-making task, involving an agent consisting of a data generation policy (which generates a plan for creating training data) and a data generation engine (which transforms the plan into data), inside an environment that provides feedback from a student. The agent’s end goal is to improve student model performance. Students are iteratively trained and evaluated on generated data, with their feedback (in the form of errors or weak skills) being reported to the agent after each iteration. As a general-purpose testbed, DataEnvGym includes multiple instantiations of teacher environments across three levels of structure in the state representation and action space, with varying levels of scaffolding support. More structured environments are based on automatically-inferred skills and offer a higher degree of interpretability and control over the curriculum. We support developing and testing data generation agents in three diverse tasks covering both text and images (mathematics, programming, and visual question answering) and test multiple student and teacher models. We find that example agents in our teaching environments can iteratively improve students across diverse tasks and settings. Moreover, we show that environments can teach different skill levels and can be used to test variants of key modules, pointing to directions of future work in improving data generation agents, engines, and feedback mechanisms. Project page: https://DataEnvGym.github.io.
@@ 1 Introduction
We propose DataEnvGym, a testbed – or gym – of parameterizable teacher environments for developing autonomous data generation agents (i.e. teachers), whose goal is to iteratively improve student models by generating targeted training data conditioned on representations of the student’s weaknesses (e.g. errors, inferred weak skills). We frame this task as an iterative interaction between a teacher agent and a student model in which the teacher agent creates training data that improves the student model. In the same way that game environments can be used to evaluate game-playing agents in conventional reinforcement learning (RL), DataEnvGym’s modular environments (cf. Sec. 2.1) allow us to test data generation agents for a given student model. In these environments, the data generation agent (teacher), performs multiple rounds of data creation and receives feedback from the student after each iteration in the form of student performance, which is the teacher’s reward signal. We provide modules for data generation, training, and evaluation, with final performance being measured by the improvement to the student model. An overview of the DataEnvGym can be seen in Fig. 1. First, the environment provides the agent with a state $s_{t}$, containing information about the errors of current student model $m_{t}$ (cf. Fig. 1(a)). Then, the agent’s data generation policy $\pi$ predicts actions that constitute a plan for generating training data (cf. Fig. 1(b)): $p_{t}\sim\pi(p_{t}|s_{t})$. Next, the agent’s data generation engine executes the plan to create training data $d_{t}$ (cf. Fig. 1(c)). The created datapoints are then used to train an updated student model (cf. Fig. 1(d)): $m_{t+1}=\texttt{Train}(m_{t},d_{t})$. The updated student model is re-evaluated to produce the next iteration’s state $s_{t+1}$ and provide feedback to the agent (in the form of student performance). DataEnvGym is designed in a generalizable way to support data creation for diverse agents across multiple tasks, covering multimodal (visual question answering) and text-only (mathematics and programming) tasks.
@@
We benchmark several baseline agents as examples (data generation policies combined with data generation engines) in DataEnvGym’s teaching environments, across different domains (mathematics, programming, visual question answering) and on different student and teacher models. Generally, we find that the example agents we provide already improve student performance when models are trained in DataEnvGym’s teaching environments; after training, students see a consistent improvement when compared to their starting point (i.e. before training in the environment). Across environments, students improve by an average of $4.43\%$ (absolute accuracy) on GQA, $4.82\%$ on MATH, and $1.80\%$ on LiveCodeBench. Moreover, we find that our example agents can make use of student feedback to help iteratively improve the student: we compare baseline agent policies that make use of student feedback states (“With State” policies) to ones that do not (“No State” policies), finding that conditioning on the feedback state is key to successful data generation, with “With State” policies outperforming “No State” by $3.5\%$ in Open-Ended, $2.05\%$ in Skill-List, and $1.08\%$ in Skill-Tree. We also show that some environments make improving students more challenging for teachers, based on how flexible versus how controllable (and interpretable) the curriculum needs to be. Moreover, we show that environments can teach different skills (based on skill frequency and question difficulty) and can be used to test variants of key modules, e.g., the skill discovery module. Lastly, we provide qualitative examples of student model predictions before and after our training. Overall, DataEnvGym is a general-purpose testbed for developing and evaluating data generation agents, engines, and feedback mechanisms, laying the foundation for future improvements to these key elements of automated model improvement.
@@ 3 Experiments
We experiment with DataEnvGym environments in three domains: visual questions answering, mathematics, and programming. For visual question answering, we use the GQA (Hudson & Manning, 2019); for mathematics, we use MATH (Hendrycks et al., 2021); for programming, we use LiveCodeBench (Jain et al., 2024). For most experiments (reported in Sec. 3.1) we start from instruction-tuned models rather than base models because we believe it is a more realistic and challenging setting since starting from instruction tuned models is standard for applications and these models have undergone post-training on large amounts of task-specific data. For GQA, we use PaliGemma-3b-pt-224 (Beyer et al., 2024) as our student model and we use GPT-4o OpenAI (2024) as the teacher agent policy, augmented with SDXL-Turbo (Sauer et al., 2023) for T2I generation. For MATH and LiveCodeBench, we use Gemma-2-2B-Instruct (Gemma Team, 2024) and Llama-3-8B-Instruct (Llama Team, 2024) as students respectively, and generate data with GPT-4o. Note that the student models we use are typically already proficient at the target task and thus are difficult to improve. For each domain, we choose the student model to satisfy the following criteria: 1) the student should be strong enough to perform the given task (e.g. LiveCodeBench is too challenging for a Gemma2-2B student). 2) The student should not have been heavily post-trained s.t. further improvements are unlikely (e.g. Llama3-8B has been extensively trained for math and further improvements with additional training are unlikely, regardless of the data generation agent). Details can be found in Sec. B.6 (validation and test splits) and Sec. B.3 (data generation).
@@ 3.1 Primary Results: VQA, Mathematics, Programming
Teaching is easier in some environments than others. With a fixed student and task (i.e. looking at “With State” entry across the columns of Tab. 2), teachers typically elicit the highest student performance in the unconstrained Open-Ended environments, where they are not required to teach a specific set of skills. However, there may be domain specific effects here as the teachers in the Skill-Tree environment perform the best on the multimodal GQA dataset (+5.58%), whereas this is reversed for MATH, where teachers in the Open-Ended environment perform the best (+7.66%).
@@
These difference may relate to the level of constraint imposed: in the Open-Ended environment, the teacher can produce any data without any constraints, while in the skill-structured environments, we require the teacher to improve the student along specified skills. This may be a more difficult task, as it may impose suboptimal constraints on the teacher, i.e., the teacher may have difficulty teaching the specified skills, whereas in the unconstrained Open-Ended environment, the teacher may be implicitly able to identify and teach the skills it is best at teaching. However, unconstrained teaching (which has no clear link between errors and generated data) may not always be useful in practice, e.g. when a user wants to control the training and make targeted improvements to certain skills.
@@
Figure 4: Per-skill accuracy improvements of Gemma-2B trained on MATH in the Skill-Tree environment, as a function of (a) question difficulty and (b) skill rarity (inverse of frequency) in the training data. The biggest performance increases occur in the middle range for difficulty and on the lower range for rarity (i.e. on more frequent skills).
@@ 3.2 Analysis: Difficulty/Rarity, Training Dynamics, Skill Discovery Quality, and Qualitative Examples
Figure 5: Training dynamics across three tasks. Performance generally increases over iterations.
Iterative training dynamics. In Fig. 5, we plot the change in the student model’s performance on the validation set throughout a full run in DataEnvGym on each task and for each environment. Each experiment is truncated once the performance consistently decreases for multiple iterations. We use the “With State” baseline agents for each environment, and use the same models as in Tab. 2. Fig. 5 shows that the students generally improve across iterations. In other words, the baseline agents do uncover new datapoints that further improve the student at each iteration. Moreover, the behavior of agents trained in the three environments differs across tasks: on GQA, Skill-Tree improves for more iterations than the other environments, while on MATH it reaches a plateau and on LiveCodeBench it is truncated after two rounds of data generation.
@@ 4 Related Work
Training Environment Generation. In agent learning frameworks, designing training environments usually becomes a bottleneck, as it requires sophisticated human efforts (Park et al., 2024). Unsupervised environment design (UED) (Dennis et al., 2020; Jiang et al., 2021; Parker-Holder et al., 2022) explores progressively increased environment difficulty based on agent scores in simple games. Liesen et al. (2024) introduce a meta-learning approach to create learning environments for continuous control. In vision-language navigation (VLN), past work (Li et al., 2022; Li & Bansal, 2024; Wang et al., 2023c) propose augmenting the visual diversity of training environments with image generation models. Generation has been applied to game environments as well: Cobbe et al. (2020) generate a diverse set of game environments for training RL agents and measuring their generalization. In our previous work – EnvGen (Zala et al., 2024) – we continuously adapt game instances for training RL agents, using an LLM to generate different game engine parameters that teach core skills to the agent based on feedback from the agents’ intermediate progress, which improves the final performance of agents as well as learning efficiency. Yang et al. (2024) generate 3D embodied environments from user-specified prompts, generating rooms in different styles. Generally, past environment generation work has focused on restricted environments like simple games, often with a few predefined actions and skills. Moreover, past environment generation work has focused on developing students rather than improving the data generation process itself. In contrast, our present work focuses on data generation agents, creating a testbed for teachers and treating students as part of the environment. Furthermore, our work introduces environments for data generation with automatic skill discovery in a diverse set of open-ended and more realistic settings such as mathematics, visual question answering, and programming.
@@
Learning from Generated Data. DataEnvGym involves transferring task knowledge from a teacher agent to a student model in an effective way, based on the student model’s feedback. Past work on knowledge transfer from one model to another has been centered around knowledge distillation, where outputs from larger models are used to train smaller ones (Hinton et al., 2015; Buciluǎ et al., 2006; Chen et al., 2020); unlike the process in DataEnvGym, this process is typically not adaptive, relying on fixed datasets of inputs that are processed by the larger teacher model and used to train the student. In the context of LLMs, symbolic distillation (West et al., 2022) has become increasingly common; here, text is generated from a large model and used to train a smaller one, e.g., in instruction tuning (Wang et al., 2023a) or distilling chain-of-thought reasoning (Wei et al., 2022) from large proprietary models into smaller models (Magister et al., 2023; Shridhar et al., 2023; Fu et al., 2023; Ho et al., 2023; Saha et al., 2023; Mukherjee et al., 2023; Mitra et al., 2023; Chen et al., 2024). The kind of teaching that data generation agents are expected to perform in DataEnvGym’s environments differs from the knowledge distillation setting in that the inputs to the model themselves are model-generated (as opposed to being sourced from an existing dataset). Moreover, the inputs are dynamically determined based on the student model’s feedback and errors, whereas in knowledge distillation they are determined by a fixed dataset or generated all at once. Note that DataEnvGym is compatible with different methods of training the student (i.e., symbolic distillation, logit-based losses, etc.), and these can be swapped into our modular environments.

### v2

@@ Abstract
The process of creating training data to teach models is currently driven by humans, who manually analyze model weaknesses and plan how to create data that improves a student model. Recent approaches using large language models (LLMs) as annotators reduce human annotation effort, but still require humans to interpret feedback from evaluations and control the LLM to produce data the student needs. Automating this labor-intensive process by creating autonomous data generation agents – or teachers – is desirable, but requires environments that can simulate the feedback-driven, iterative, closed loop of data creation. To enable rapid and scalable testing for such agents and their modules, we introduce DataEnvGym, a testbed of teacher environments for data generation agents. DataEnvGym frames data generation as a sequential decision-making task, involving an agent consisting of a data generation policy (which generates a plan for creating training data) and a data generation engine (which transforms the plan into data), inside an environment that provides feedback from a student. The agent’s end goal is to improve student model performance. Students are iteratively trained and evaluated on generated data, with their feedback (in the form of errors or weak skills) being reported to the agent after each iteration. As a general-purpose testbed, DataEnvGym includes multiple instantiations of teacher environments across three levels of structure in the state representation and action space, with varying levels of scaffolding support. More structured environments are based on automatically-inferred skills and offer a higher degree of interpretability and control over the curriculum. We support developing and testing data generation agents in four diverse tasks covering text, images, and actions (mathematics, programming, visual question answering, and tool-use) and test multiple student and teacher models. We find that example agents in our teaching environments can iteratively improve students across diverse tasks and settings. Moreover, we show that environments can teach different skill levels and can be used to test variants of key modules, pointing to directions of future work in improving data generation agents, engines, and feedback mechanisms. Project page: https://DataEnvGym.github.io.
@@ 1 Introduction
We propose DataEnvGym, a testbed – or gym – of parameterizable teacher environments for developing autonomous data generation agents (i.e., teachers), whose goal is to iteratively improve student models by generating targeted training data conditioned on representations of the student’s weaknesses (e.g,. errors, inferred weak skills). We frame this task as an iterative interaction between a teacher agent and a student model in which the teacher agent creates training data that improves the student model. In the same way, that game environments can be used to evaluate game-playing agents in conventional reinforcement learning (RL), DataEnvGym’s modular environments (cf. Sec. 2.1) allow us to test data generation agents for a given student model. In these environments, the data generation agent (teacher), performs multiple rounds of data creation and receives feedback from the student after each iteration in the form of student performance, which is the teacher’s reward signal. We provide modules for data generation, training, and evaluation, with final performance being measured by the improvement to the student model. An overview of the DataEnvGym can be seen in Fig. 1. First, the environment provides the agent with a state $s_{t}$, containing information about the errors of current student model $m_{t}$ (cf. Fig. 1(a)). Then, the agent’s data generation policy $\pi$ predicts actions that constitute a plan for generating training data (cf. Fig. 1(b)): $p_{t}\sim\pi(p_{t}|s_{t})$. Next, the agent’s data generation engine executes the plan to create training data $d_{t}$ (cf. Fig. 1(c)). The created datapoints are then used to train an updated student model (cf. Fig. 1(d)): $m_{t+1}=\texttt{Train}(m_{t},d_{t})$. The updated student model is re-evaluated to produce the next iteration’s state $s_{t+1}$ and provide feedback to the agent (in the form of student performance). DataEnvGym is designed in a generalizable way to support data creation for diverse agents across multiple tasks, covering multimodal (visual question answering) and text-only (mathematics and programming) tasks.
@@
We benchmark several baseline agents as examples (data generation policies combined with data generation engines) in DataEnvGym’s teaching environments, across different domains (mathematics, programming, visual question answering) and on different student and teacher models. Generally, we find that the example agents we provide already improve student performance when models are trained in DataEnvGym’s teaching environments; after training, students see a consistent improvement when compared to their starting point (i.e., before training in the environment). Across environments, students improve by an average of $4.43\%$ (absolute accuracy) on GQA, $4.82\%$ on MATH, $1.80\%$ on LiveCodeBench, $15.17\%$ on NaturalBench, and $20.81\%$ on MnMs. Moreover, we find that our example agents can make use of student feedback to help iteratively improve the student: we compare baseline agent policies that make use of student feedback states (“With State” policies) to ones that do not (“No State” policies), finding that conditioning on the feedback state is key to successful data generation, with “With State” policies outperforming “No State” by $3.5\%$ in Open-Ended, $2.05\%$ in Skill-List, and $1.08\%$ in Skill-Tree. We also show that some environments make improving students more challenging for teachers, based on how flexible versus how controllable (and interpretable) the curriculum needs to be. Moreover, we show that environments can teach different skills (based on skill frequency and question difficulty) and can be used to test variants of key modules, e.g., the skill discovery module. Lastly, we provide qualitative examples of student model predictions before and after our training. Overall, DataEnvGym is a general-purpose testbed for developing and evaluating data generation agents, engines, and feedback mechanisms, laying the foundation for future improvements to these key elements of automated model improvement.
@@ 3 Experiments
We experiment with DataEnvGym environments in four domains: visual question answering, mathematics, programming and tool-use. For visual question answering, we use GQA (Hudson & Manning, 2019) and NaturalBench (Li et al., 2024); for mathematics, we use MATH (Hendrycks et al., 2021); for programming, we use LiveCodeBench (Jain et al., 2024); for tool-use, we use MnMs (Ma et al., 2024). For most experiments (reported in Sec. 3.1), we start from instruction-tuned models rather than base models because we believe it is a more realistic and challenging setting since starting from instruction-tuned models is standard for applications, and these models have undergone post-training on large amounts of task-specific data. For GQA and NaturalBench, we use PaliGemma-3b-pt-224 (Beyer et al., 2024) as our student model and we use GPT-4o OpenAI (2024) as the teacher agent policy, augmented with SDXL-Turbo (Sauer et al., 2023) for T2I generation. For MATH, we use Gemma-2-2B-Instruct (Gemma Team, 2024) as a student. For LiveCodeBench and MnMs, we use Llama-3-8B-Instruct (Llama Team, 2024) as a student. For all domains, we generate data with GPT-4o. Note that the student models we use are typically already proficient at the target task and thus are difficult to improve. For each domain, we choose the student model to satisfy the following criteria: 1) the student should be strong enough to perform the given task (e.g., LiveCodeBench is too challenging for a Gemma2-2B student). 2) The student should not have been heavily post-trained s.t. further improvements are unlikely (e.g., Llama3-8B has been extensively trained for math and further improvements with additional training are unlikely, regardless of the data generation agent). Details can be found in Sec. B.7 (validation and test splits) and Sec. B.3 (data generation). Experiments on tool-use (MnMs) and additional multimodal experiments (NaturalBench) are detailed in Sec. 3.2.
@@ 3.1 Primary Results: VQA, Mathematics, Programming
Teaching is easier in some environments than others. With a fixed student and task (i.e., looking at “With State” entry across the columns of Tab. 2), teachers typically elicit the highest student performance in the unconstrained Open-Ended environments, where they are not required to teach a specific set of skills. However, there may be domain-specific effects here as the teachers in the Skill-Tree environment perform the best on the multimodal GQA dataset (+5.58%), whereas this is reversed for MATH, where teachers in the Open-Ended environment perform the best (+7.66%).
@@
These differences may relate to the level of constraint imposed: in the Open-Ended environment, the teacher can produce any data without any constraints, while in the skill-structured environments, we require the teacher to improve the student along specified skills. This may be a more difficult task, as it may impose suboptimal constraints on the teacher, i.e., the teacher may have difficulty teaching the specified skills, whereas in the unconstrained Open-Ended environment, the teacher may be implicitly able to identify and teach the skills it is best at teaching. However, unconstrained teaching (which has no clear link between errors and generated data) may not always be useful in practice, e.g., when a user wants to control the training and make targeted improvements to certain skills.
3.2 Additional Multimodal and Tool-Use Benchmarks
In this section, we show results on additional challenging multimodal and tool-use tasks: NaturalBench (Li et al., 2024) and Table 3 and MnMs (Ma et al., 2024) in Table 4. MnMs and NaturalBench include fine-grained metrics other than accuracy, which enables a more detailed view of how data generated by different methods affects a student model. On both MnMs and NaturalBench, conditioning on the task and errors of the student allows the teacher to generate more effective training data (compare “With State” against “No State” in Table 3 and Table 4).
On NaturalBench, the training data generated when conditioning on the student’s state is more effective at improving the student on all metrics. The “Group Score” probes the consistency of a model across four visual question-answer pairs that test the same concept. The “Group Score” increase from generated data without conditioning on the state is zero, while conditioning on the state results in an increase in the group score.
MnMs is a tool-use benchmark that requires a modified skill discovery mechanism, as the standard pipeline applied to MnMs produces skills that are not tightly coupled to the underlying tools required. For MnMs, we consider each combination of tools as a “skill”. This results in a very large number of “skills” for tool-use. To generate data on a fixed budget, we must sample a subset of skills (tool combinations) for which to generate data. We experiment with several sampling methods, and find that the most effective sampling method is to randomly sample a subset of skills for which the student does not have near-perfect accuracy.
We refer the reader to the GitHub repository for complete details of prompts and hyperparameters.
Table 3: Performance metrics for NaturalBench (Li et al., 2024) before and after teaching using the Open-Ended and Skill-List environments.
Question Score Image Score Binary Score Group Score Avg. Improvement
Before teaching 3.92 3.92 48.53 0.00 -
Open-Ended environment
+With State $\pi(p_{t}|s_{t})$ 26.00 (+22.08%) 29.00 (+25.08%) 61.00 (+12.47%) 6.00 (+6.00%) (+16.41%)
+No State $\pi(p_{t}|\cdot)$ 11.00 (+7.08%) 9.00 (+5.08%) 51.50 (+2.97%) 0.00 (+0.00%) (+3.78%)
Skill-List environment
+With State $\pi(p_{t}|s_{t})$ 25.00 (+21.08%) 26.00 (+22.08%) 59.50 (+10.97%) 2.00 (+2.00%) (+14.03%)
+No State $\pi(p_{t}|\cdot)$ 11.00 (+7.08%) 9.00 (+5.08%) 51.50 (+2.97%) 0.00 (+0.00%) (+3.78%)
Table 4: Performance metrics for MnMs (Ma et al., 2024) before and after teaching using the Open-Ended and Skill-List environments.
Plan Accuracy Tool Precision Tool Recall Tool F1 Avg. Gain
Before teaching 37.30 73.74 68.18 65.41 -
Open-Ended environment
+With State $\pi(p_{t}|s_{t})$ 72.00 (+34.70%) 87.32 (+13.98%) 82.62 (+14.44%) 82.25 (+16.84%) (+19.99%)
+No State $\pi(p_{t}|\cdot)$ 62.13 (+24.83%) 87.94 (+14.60%) 87.09 (+12.91%) 80.51 (+15.10%) (+16.86%)
Skill-List environment
+With State $\pi(p_{t}|s_{t})$ 75.96 (+38.66%) 86.11 (+12.77%) 84.35 (+16.17%) 84.36 (+18.95%) (+21.64%)
+No State $\pi(p_{t}|\cdot)$ 62.13 (+24.83%) 87.94 (+14.60%) 87.09 (+12.91%) 80.51 (+15.10%) (+16.86%)
@@ 3.2 Additional Multimodal and Tool-Use Benchmarks
Figure 4: Per-skill accuracy improvements of Gemma-2B trained on MATH in the Skill-Tree environment, as a function of (a) question difficulty and (b) skill rarity (inverse of frequency) in the training data. The biggest performance increases occur in the middle range for difficulty and in the lower range for rarity (i.e., on more frequent skills).
@@ 3.3 Analysis: Difficulty/Rarity, Training Dynamics, Skill Discovery Quality, and Qualitative Examples
Figure 5: Training dynamics across three tasks. Each line is split into a solid segment terminating in a vertical line (the maximum performance achieved by the student) and a dashed line showing the effect of continued training beyond this point.
Iterative training dynamics. In Fig. 5, we plot the change in the student model’s performance on the validation set throughout a full run in DataEnvGym on each task and for each environment. Each experiment is truncated once the performance saturates (consistently fails to increase for multiple iterations). We use the “With State” baseline agents for each environment, and use the same models as in Tab. 2. Fig. 5 shows that the students generally improve across iterations. In other words, the baseline agents do uncover new datapoints that further improve the student at each iteration. Moreover, the behavior of agents trained in the three environments differs across tasks: on GQA, Skill-Tree improves for more iterations than the other environments, while on MATH it reaches a plateau and on LiveCodeBench it is truncated after two rounds of data generation.
@@ 4 Related Work
Training Environment Generation. In agent learning frameworks, designing training environments usually becomes a bottleneck, as it requires sophisticated human efforts (Park et al., 2024). Unsupervised environment design (UED) (Dennis et al., 2020; Jiang et al., 2021; Parker-Holder et al., 2022) explores progressively increased environment difficulty based on agent scores in simple games and Ruiz et al. (2019) control the distribution of synthesized data in order to maximize the accuracy of a model trained on that data. Liesen et al. (2024) introduce a meta-learning approach to create learning environments for continuous control. In vision-language navigation (VLN), past work (Li et al., 2022; Li & Bansal, 2024; Wang et al., 2023c) propose augmenting the visual diversity of training environments with image generation models. Generation has been applied to game environments as well: Cobbe et al. (2020) generate a diverse set of game environments for training RL agents and measuring their generalization. In our previous work – EnvGen (Zala et al., 2024) – we continuously adapt game instances for training RL agents, using an LLM to generate different game engine parameters that teach core skills to the agent based on feedback from the agents’ intermediate progress, which improves the final performance of agents as well as learning efficiency. Yang et al. (2024) generate 3D embodied environments from user-specified prompts, generating rooms in different styles. Sudhakaran et al. (2023) and Todd et al. (2023) use LLMs for open-ended generation of game levels for tile-based games. OMNI (Zhang et al., 2023) uses LLMs for task selection in open-ended environments, while OMNI-Epic (Faldor et al., 2024) augments OMNI with the ability to generate new tasks in simulated robotics settings. Generally, past environment generation work has focused on restricted environments like simple games, often with a few predefined actions and skills. Moreover, past environment generation work has focused on developing students rather than improving the data generation process itself. In contrast, our present work focuses on data generation agents, creating a testbed for teachers and treating students as part of the environment. Furthermore, our work introduces environments for data generation with automatic skill discovery in a diverse set of open-ended and more realistic settings such as mathematics, visual question answering, and programming.
@@
Learning from Generated Data. DataEnvGym involves transferring task knowledge from a teacher agent to a student model in an effective way, based on the student model’s feedback. Past work on knowledge transfer from one model to another has been centered around knowledge distillation, where outputs from larger models are used to train smaller ones (Hinton et al., 2015; Buciluǎ et al., 2006; Chen et al., 2020); unlike the process in DataEnvGym, this process is typically not adaptive, relying on fixed datasets of inputs that are processed by the larger teacher model and used to train the student. In the context of LLMs, symbolic distillation (West et al., 2022) has become increasingly common; here, text is generated from a large model and used to train a smaller one, e.g., in instruction tuning (Wang et al., 2023a) or distilling chain-of-thought reasoning (Wei et al., 2022) from large proprietary models into smaller models (Magister et al., 2023; Shridhar et al., 2023; Fu et al., 2023; Ho et al., 2023; Saha et al., 2023; Mukherjee et al., 2023; Mitra et al., 2023; Chen et al., 2024). The kind of teaching that data generation agents are expected to perform in DataEnvGym’s environments differs from the knowledge distillation setting in that the inputs to the model themselves are model-generated (as opposed to being sourced from an existing dataset). Moreover, the inputs are dynamically determined based on the student model’s feedback and errors, whereas in knowledge distillation, they are determined by a fixed dataset or generated all at once. Note that DataEnvGym is compatible with different methods of training the student (i.e., symbolic distillation, logit-based losses, etc.), and these can be swapped into our modular environments.